OpenStudy (anonymous):

the value of c for which f(x)=x+(c/x) has a local minimum at x=3 is?

zepdrix (zepdrix):

If x=3 is a local minimum point, then x=3 corresponds to a critical point, ya?\[\large\rm f'(3)=0\]

zepdrix (zepdrix):

Take derivative, set equal to 0, (since we're looking for critical points) Plug in x=3, (since that corresponds to a critical point) solve for c.

OpenStudy (anonymous):

so x+c/x is 1+c(x^-1) right then c-x^-2=0?

zepdrix (zepdrix):

woah not sure what happened there, looks like you took derivative of the first term twice or something.

zepdrix (zepdrix):

and yes, the negative comes down, but make sure you're not wring c MINUS x^-2, that's multiplication

zepdrix (zepdrix):

\[\large\rm x+cx^{-1}\qquad\to\qquad 1+-cx^{-2}\]

zepdrix (zepdrix):

writing* not wring lol

OpenStudy (anonymous):

oh okay. 1-cx^-2=0 -1/-(3^-2)=9 Thanks!

zepdrix (zepdrix):

9? Good good good :)